((a+b)^2)-((a-c)^2) solve to the least common factor.
Answer by mathchemprofessor(65)

(a+b)^2)-(a-c)^2)
The above expression is of the form,x^2-y^2=(x+y)*(x-y)
So,
(a+b)^2)-(a-c)^2)=((a+b)+(a-c))*((a+b)-(a-c))=(a+b+a-c)*(a+b-a+c)
=(2a+b-c)*(b+c)